OpenStudy (anonymous):

y= 2/5x-3 in standard form

jimthompson5910 (jim_thompson5910):

First multiply both sides by 5 to clear out the fraction y= 2/5x-3 5y= 5*(2/5x-3) 5y = 2x - 15

jimthompson5910 (jim_thompson5910):

Then you need to get the equation into Ax + By = C form, so you subtract 2x from both sides 5y = 2x - 15 5y-2x = -15 -2x + 5y = -15 and finally, multiply both sides by -1 to make the x coefficient positive -2x + 5y = -15 -1(-2x + 5y) = -1*(-15) 2x - 5y = 15

OpenStudy (anonymous):

thanks